OpenStudy (anonymous):

How would you write log627 – 2log63 as a single logarithm?

OpenStudy (anonymous):

log627- 2log63

OpenStudy (michele_laino):

please you have to apply this properties: \[n*\log _{b}x=\log _{b}(x ^{n})\] and subsequently: \[\log _{b}x-\log _{b}y=\log _{b}(\frac{ x }{ y })\]

OpenStudy (michele_laino):

for example, we have \[2\log 63=\log(63^{2})=...\] please continue

OpenStudy (triciaal):

when you subtract logs you divide the numbers 27 can be expressed as 3 * 9 and 63 can be expressed as 7 * 9 when you multiply numbers add the logs also use the log rule above for exponents. oops that is not log 63 its log base 6 of 3 so log base 6 of 3^2

OpenStudy (triciaal):

27 = 3^3